OpenStudy (anonymous):

f (x) = 3x + 6, g(x) = 2x2 Find (fg)(x

OpenStudy (anonymous):

Any idea

OpenStudy (anonymous):

I think its 6x^2 +6

OpenStudy (anonymous):

Thanks for the free medal!

OpenStudy (anonymous):

And its wrong @Faith_123

OpenStudy (anonymous):

Then is it 6x^3 +12×^2

OpenStudy (anonymous):

Wrong!!

OpenStudy (anonymous):

How do I solve this!?

OpenStudy (anonymous):

Not fair U took my medal!

OpenStudy (anonymous):

6x^2 +6 is correct! I was kidding

OpenStudy (jdoe0001):

\(\bf f (x) = 3x + 6\qquad g(x) = 2x^2 \\ \quad \\ (fg)(x)\implies f(x)g(x)\implies f(x)\times g(x)\implies (3x+6)\times (2x^2)\)

OpenStudy (anonymous):

(Fg)(x)

OpenStudy (anonymous):

It's a multiple choice question and that's not one of the options @ KissMyAxe

OpenStudy (jdoe0001):

so what do you get if you multiply them over?

OpenStudy (anonymous):

First find `(fg)`which is `6x^2+ 6`

OpenStudy (anonymous):

than multiply this by 2x^2

OpenStudy (anonymous):

Did u get it?

OpenStudy (anonymous):

Not really...

OpenStudy (anonymous):

(6x^2+ 6)*2x^2, now

OpenStudy (anonymous):

Yes but that gives you 12x^4 + 12x^2

OpenStudy (anonymous):

Yeah

OpenStudy (anonymous):

Why u typing so slow

OpenStudy (anonymous):

That is not an option to this multiple choice question

OpenStudy (jdoe0001):

hmmm

OpenStudy (anonymous):

No idea than!

OpenStudy (jdoe0001):

\(\large \bf 6x^3 +12x^2\) is correct for (fg)(x)

OpenStudy (jdoe0001):

unless you're expected to factor or such

OpenStudy (anonymous):

How did you get that answer?

OpenStudy (jdoe0001):

hmm? you got that answer yourself " Then is it 6x^3 +12×^2 " <----

OpenStudy (jdoe0001):

so..how did You get it then?

OpenStudy (anonymous):

Where did i go wrong?

OpenStudy (jdoe0001):

\(\bf f (x) = 3x + 6\qquad g(x) = 2x^2 \\ \quad \\ (fg)(x)\implies f(x)g(x)\implies f(x)\times g(x)\implies (3x+6)\times (2x^2) \\ \quad \\ 6x^3+12x^2\implies6x^2(x+2)\)

OpenStudy (anonymous):

OK that makes sense thanks